# No more hard prompts: SoftSRV prompting for synthetic data generation

### v2

@@ Abstract
We present a novel soft prompt based framework, SoftSRV, that leverages a frozen pre-trained large language model (LLM) to generate targeted synthetic text sequences. Given a sample from the target distribution, our proposed framework uses data-driven loss minimization to train a parameterized “contextual” soft prompt. This soft prompt is then used to steer the frozen LLM to generate synthetic sequences that are similar to the target distribution. We argue that SoftSRV provides a practical improvement over common hard-prompting approaches that rely on human-curated prompt-templates, which can be idiosyncratic, labor-intensive to craft, and may need to be specialized per domain. We empirically evaluate SoftSRV and hard-prompting baselines by generating synthetic data to fine-tune a small Gemma model on three different domains (coding, math, reasoning). To stress the generality of SoftSRV, we perform these evaluations without any particular specialization of the framework to each domain. We find that SoftSRV significantly improves upon hard-prompting baselines, generating data with superior fine-tuning performance and that better matches the target distribution according to the mauve similarity metric.
@@ 1 Introduction
In recent years, pre-trained large language models (LLMs) have proven to be effective in generating synthetic natural language training data (Gunasekar et al., 2023; Li et al., 2023; Eldan & Li, 2023; Mukherjee et al., 2023; Mitra et al., 2023; Abdin et al., 2024). This is particularly true when the synthetic data is used to pre-train or fine-tune smaller language models, enabling performances that rival models that are orders of magnitude larger (Liu et al., 2023). There are several motivations for generating and using synthetic training data; chief among them is the need to train models for domains where little natural high-quality text may be readily available or may be difficult to procure.
In order to generate synthetic text, a significant amount of human-driven prompt engineering is invested into developing prompts that steer the generating LLM into producing high-quality text from a targeted domain while also encouraging sufficient diversity. This point was very nicely summed up by the authors of the open-source synthetic text repository Cosmopedia (Ben Allal et al., 2024), when recounting their attempt to recreate a large synthetic dataset similar to the one generated to train Phi 1.5 (Li et al., 2023):
“Heads up: If you are anticipating tales about deploying large-scale generation tasks across hundreds of H100 GPUs, in reality most of the time for Cosmopedia was spent on meticulous prompt engineering.”  – Ben Allal et al. (2024)
@@
To address these issues, we propose an algorithmic framework, Soft prompt-based Synthesis with Randomized Variation (SoftSRV), that leverages soft-prompting (also known as prompt-tuning) for synthetic text data generation. Soft prompt training is a parameter efficient tuning method and requires a relatively limited amount of compute (Lester et al., 2021). Perhaps equally important, since the soft prompt is trained using a data-driven training algorithm, it requires essentially no human-in-the-loop intervention, enabling the process to be readily deployed across many different domains. Furthermore, soft prompts themselves can allow for more expressive input contexts to the generating LLM compared to natural language hard prompts. A soft “token”, represented by a dense vector, is not restricted to correspond to a particular discrete natural language token (e.g., sub-word or character). This intuitive observation is formalized in Petrov et al. (2024b), which shows that in specific settings soft prompts can induce an LMM to produce an exponential (in sequence length) number of text completions, while hard prompts only allow for a linear number of completions.
Why do we expect soft-prompting to be effective for targeted synthetic data generation? Prior theoretical research on fine-tuning language models suggests that a data-driven optimization of soft prompts guides a pre-trained model towards specific concepts or tasks it has already learned, essentially steering the model towards a relevant subspace of interest (Wies et al., 2023; Petrov et al., 2024b; a). Our goal is to use soft prompts to steer the pre-trained model towards generating text that most resembles the target distribution. Subsequently, fine-tuning a smaller model using the generated data provides an effective way to transfer knowledge from the larger model to the smaller model.
Unlike in typical prompt-tuning approaches, we do not prepend a soft prompt to an existing hard prompt, but instead use the soft prompt alone as input context to the LLM. We use a sample of text sequences (i.e., a sample from the desired target text distribution) and language-modeling loss to learn a parametrized soft prompt. Once trained, the soft prompt can be varied by conditioning on a context vector derived from an example sequence, allowing for additional expressive power and the potential to better fit different regions of a potentially complex target distribution.
Our contributions presented in this work are as follows:
  • We demonstrate that soft prompts can be effectively trained for the purpose of targeted synthetic text generation used to fine-tune downstream models.
  • We investigate the value of learning parameterized families of soft prompts that can be conditioned on an input context, finding they are critical for best fitting complex target distributions.
  • Our empirical evaluations on coding, math, and reasoning tasks find superior downstream performance for models fine-tuned on SoftSRV generated text compared to that of models fine-tuned on data generated by baseline hard-prompting approaches.
  • We show that the SoftSRV approach is general and offers greater versatility than hard-prompting approaches as it can be readily applied across different domains with minimal manual intervention.
@@ 2 Proposed Approach
In this section, we introduce the general SoftSRV framework as well as a few specific instantiations that are studied in this work. First, we start with some basic notation and terminology.
Given a vocabulary $\mathcal{V}$ of textual tokens, let $\{x_{1},\ldots,x_{n}\}$ denote a sample of $n$ text sequences, belonging to the set of all possible sequences $S^{m}$ of a finite maximum length $m$, drawn according to a fixed but unknown distribution $\mathcal{D}$. Although we are not able to directly sample additional sequences from $\mathcal{D}$, our goal is to synthesize sequences that could have plausibly been drawn according to $\mathcal{D}$. We assume access to a (frozen) LLM, denoted $L:S^{m}\to S^{m}$, where we input and output sequences of equal fixed length $m$ for notational simplicity and without loss of generality. Furthermore, we explicitly decompose the LLM, $L=H\circ E$, where $E:S^{m}\to\mathbb{R}^{d\times m}$ represents the initial embedding layer that embeds each token of the input sequence to a $d$-dimensional dense vector, and $H:\mathbb{R}^{d\times m}\to S^{m}$ represent the remainder of the language model that maps the embedded tokens to the output sequence. In contrast to the prompt tuning methods of Lester et al. (2021) in the standard fine-tuning setting, we do not prepend the learned soft tokens to a hard prompt after it passes the initial embedding layer, $E$. Instead, for our approach we discard $E$ entirely method and rely on the frozen model $H$.
The SoftSRV framework seeks to synthesize sequences similar to those drawn from $\mathcal{D}$ by training a “soft prompt”, i.e. a dense embedding (or parameterized family of embeddings) $\mathbf{P}\in\mathbb{R}^{d\times t}$ consisting of $0<t<m$ “soft-tokens”. A successfully trained soft prompt, $\mathbf{P}$, should generate a sequence $x=H(\mathbf{P})$, via frozen model $H$, that has a high likelihood of occurring under the distribution $\mathcal{D}$. More generally, we can sample several different sequences from a fixed prompt, $x,x^{\prime},x^{\prime\prime}\ldots\sim H(\mathbf{P})$, by using randomized temperature-based decoding. Although temperature sampling alone does allow for some variability, we can further increase the variety of generated text by using a contextual soft prompt, $\mathbf{P}(\cdot):\mathbb{R}^{d_{e}}\to\mathbb{R}^{d\times t}$. A contextual soft prompt can be conditioned with different context vectors $\mathbf{z}\in\mathbb{R}^{d_{e}}$, during training and generation, to induce variations of the soft prompt.
@@
Figure 1: A diagram illustrating the training workflow of the SoftSRV framework. An example sequence $x$ is embedded into a dense vector $\mathbf{z}$ via a (frozen) sequence encoder model. The SoftSRV model, parameterized by $\theta$ and conditioned on the embedding $\mathbf{z}$, produces a soft prompt $\mathbf{P}_{\theta}(\mathbf{z})$. This is then fed to a (frozen) pre-trained LLM, which produces a synthetic example $x^{\prime}$. Similar to autoencoder-based training, the gradient of a next-word-prediction “reconstruction” loss is computed and used to update the SoftSRV parameters.
@@
Before introducing specific soft prompt instantiations, we describe the SoftSRV training procedure which is common throughout and also illustrated in Figure 1. In addition to the sample of data ($x_{1},\ldots,x_{n}$) and frozen LLM ($H$), we assume access to a sequence embedding function $\mathrm{emb}(\cdot):S^{m}\to\mathbb{R}^{d_{e}}$ and we let $\theta$ denote the trainable parameters of the (contextual) soft prompt $\mathbf{P}_{\theta}(\cdot)$. During training, each training sequence is embedded $\mathbf{z}_{i}=\mathrm{emb}(x_{i})$ and used to generate a conditioned soft prompt $\mathbf{P}_{\theta}(\mathbf{z})$, which is fed into the frozen LLM $H$ to produce a new sequence $x_{i}^{\prime}\sim H(\mathbf{P}_{\theta}(\mathbf{z}_{i}))$ using the standard sequential (next token) generation. A standard causal (next-word) prediction loss, denoted $\ell(\cdot,\cdot)$, is backpropagated through the network up to the soft prompt layer $\mathbf{P}_{\theta}$, and an SGD-style update is applied to $\theta$ using the gradient $\nabla_{\theta}\ell(x_{i},x_{i}^{\prime})$. This loss can be thought of as a “reconstruction” error and the entire pipeline is akin to an auto-encoder. Viewing the pipeline through this lens, it is apparent that the sequence embedder $\mathrm{emb}(\cdot)$ should be sufficiently “lossy” in order to avoid making the learning problem trivial. This lossiness can be enforced by restricting the dimension $d_{e}$ of the embedding, for example.
@@
Once the contextual soft prompt $\mathbf{P}(\mathbf{z})$ has been trained, we can then generate synthetic data by prompting the LLM using $\mathbf{P}(\mathbf{z})$ as embedded input context for different choices of context vector $\mathbf{z}$. A natural choice is to sample embeddings $(\mathbf{z}_{1},\ldots,\mathbf{z}_{n})$ derived from the data sample set $(x_{1},\ldots,x_{n})$. We now introduce a few specific SoftSRV parameterizations studied in this work.
2.1 Non-contextual soft prompt ($\mathrm{SS}_{\mathrm{NSP}}$)
@@
The simplest parameterization treats the $dt$ entries of a soft prompt, $\mathbf{P}\in\mathbb{R}^{d\times t}$, directly as trainable parameters, i.e., $\theta=\mathbf{P}$ , resulting in the following objective:
$\mathrm{argmin}_{\theta}\sum_{i=1}^{n}\ell(H(\mathbf{P}),x_{i})\,\,,$ (1)
@@
where it is understood that, in practice, the argmin over $\theta$ is only approximated via SGD. This parameterization is an instance of a non-contextual soft prompt, i.e., any context $\mathbf{z}$ is ignored. Despite the lack of context, the synthesized output may still be diversified by using non-greedy (i.e., temperature sampling) decoding during LLM generation.
2.2 Mixture of prompts ($\mathrm{SS}_{\mathrm{MPk}}$)
@@
Here, we train $k$ “basis” soft prompt matrices and define the final soft prompt, as a mixture of these bases. More precisely, in this variant the parameter set is $\theta=\{\mathbf{P}_{1},\ldots,\mathbf{P}_{k},\phi\}$, where $\mathbf{P}_{i}\in\mathbb{R}^{d\times t}$ are the basis prompts,
@@
and $W_{\phi}(\cdot):\mathbb{R}^{d_{e}}\to\mathbb{R}^{k}$ is a learned softmax function with parameters $\phi\in\mathbb{R}^{d_{w}}$. The trained $\mathrm{SS}_{\mathrm{MPk}}$ prompt is then the SGD solution to $\mathrm{argmin}_{\theta}\sum_{i=1}^{n}\ell(H(\mathbf{P}_{\theta}(\mathrm{emb}(x_{i})),x_{i})$.
The intuition behind this formulation is for each learned basis soft prompt $\mathbf{P}_{i}$ to encode a different aspect (mode) of the target data distribution and have each training example $x_{i}$ approximated by a mixture of these modes (similar to the intuition behind mixture or topic models (Hand, 2018)). Previous prompt-tuning works have also made use of a mixture of soft prompts, albeit not focused on training data synthesis (Qin & Eisner, 2021; Dun et al., 2023).
2.3 MLP Concatenated ($\mathrm{SS}_{\mathrm{MC}}$)
We consider a collection of $t$ small MLPs, whose output is concatenated to generate the final soft prompt. Let $F_{\phi_{i}}:\mathbb{R}^{d_{e}}\to\mathbb{R}^{t}$ denote the $i$th MLP with parameters $\phi_{i}$, and $\theta=\{\phi_{1},\ldots,\phi_{t}\}$ denote the parameters for the collection of MLPs. Then, we define:
@@ 2.3 MLP Concatenated ($\mathrm{SS}_{\mathrm{MC}}$)
and the trained $\mathrm{SS}_{\mathrm{MC}}$ soft prompt is the SGD solution to $\mathrm{argmin}_{\theta}\sum_{i=1}^{n}\ell(H(\mathbf{P}_{\theta}(\mathrm{emb}(x_{i})),x_{i})$.
This parameterization is the most expressive that we consider, in that each “soft-token” in the soft prompt is computed using a distinct non-linear transformation of the context vector $\mathbf{z}$.

### v3

# SoftSRV: Learn to Generate Targeted Synthetic Data

@@ Abstract
We present a novel framework, SoftSRV, that is used to generate targeted synthetic fine-tuning data for improving task-specific model performance. Given a sample from a target distribution, our proposed framework uses a data-driven loss minimization approach to steer a frozen large language model (LLM) to generate synthetic sequences that are similar to those from the target distribution. SoftSRV provides a practical improvement over common prompt engineering approaches that rely on human-engineered prompt-templates, which can be idiosyncratic, labor-intensive to craft, and may need to be specialized per domain. We empirically evaluate our method against standard baselines guiding a large LLM to generate synthetic data to fine-tune a smaller language model on three different domains (coding, math, reasoning). We perform these evaluations without any particular specialization of the framework to each domain, emphasizing the generality of our approach. We find that SoftSRV improves upon typical prompt engineering approaches, generating targeted data that leads to fine-tuned models with significantly better task-specific performance. In addition, SoftSRV-generated data better matches the target distribution according to the mauve similarity metric.
@@ 1 Introduction
In recent years, pre-trained large language models have proven to be effective in generating synthetic natural language training data (Gunasekar et al., 2023; Li et al., 2023; Eldan & Li, 2023; Mukherjee et al., 2023; Mitra et al., 2023; Abdin et al., 2024). This is particularly true when the synthetic data is used to pre-train or fine-tune smaller language models, enabling performances that rival models that are orders of magnitude larger (Liu et al., 2023). There are several motivations for generating and using synthetic training data; chief among them is the need to train models for domains where little natural high-quality text may be readily available or may be difficult to procure.
In order to generate synthetic text, a significant amount of human-driven prompt engineering is invested into developing prompts that steer the generating LLM into producing high-quality text from a targeted domain while also encouraging sufficient diversity. This point was nicely summed up by the authors of the open-source synthetic text repository Cosmopedia (Ben Allal et al., 2024), when recounting their work to recreate a large synthetic dataset similar to the one generated to train Phi 1.5 (Li et al., 2023):
“Heads up: If you are anticipating tales about deploying large-scale generation tasks across hundreds of H100 GPUs, in reality most of the time for Cosmopedia was spent on meticulous prompt engineering.”   – (Ben Allal et al., 2024)
@@
To address these issues, we propose an algorithmic framework, SoftSRV, that leverages trainable parameteric embeddings, rather than natural language prompts, to steer a pre-trained model towards generating text that most resembles the target distribution. These parametric embeddings are trained by minimizing a data-driven loss function using an autoencoder-like compression and reconstruction procedure. We restrict SoftSRV to train on only a small number of parameters, thereby using a relatively limited amount of compute. Additionally, SoftSRV requires essentially no human-in-the-loop prompt engineering, enabling the process to be readily deployed across many different domains.
Parameteric embeddings allow for more expressive inputs to an LLM compared to natural language prompts since an embedding is not restricted to correspond to a particular sequence of discrete natural language tokens. This intuitive observation is formalized in Petrov et al. (2024), which shows that in specific settings trained embeddings can induce an LMM to produce an exponential (in sequence length) number of text completions, while natural language prompts only allow for a linear number of completions. Although SoftSRV has some similarities to traditional prompt-tuning approaches, there are also several critical differences, which we discuss later in Section 2.4.
The specific contributions presented in this work are:
  • We introduce the SoftSRV framework and provide several instances of parameterized contextual embeddings that can be leveraged within the framework.
  • We demonstrate that these embeddings can be successfully trained and used to generate targeted synthetic text for fine-tuning task-specific downstream models.
  • Our empirical evaluations demonstrate that models fine-tuned on SoftSRV-generated text admit superior performance compared to those fine-tuned on data generated by baseline prompt engineering approaches or on limited amount of non-synthetic data. We show results on coding, math, and reasoning benchmarks, using both in-domain and out-of-domain tasks to study generalization of the synthetic data.
@@ 2 Proposed Approach
Figure 1: A diagram illustrating the training workflow of the SoftSRV framework. An example sequence $x$ is embedded into a dense vector $\mathbf{z}$ via a (frozen) sequence encoder model. The SoftSRV model, parameterized by $\theta$ and conditioned on the embedding $\mathbf{z}$, produces a SoftSRV embedding $\mathbf{P}_{\theta}(\mathbf{z})$. This is then fed to a (frozen) pre-trained LLM, which produces a synthetic example $x^{\prime}$. Similar to autoencoder-based training, the gradient of a next-word-prediction “reconstruction” loss is computed and used to update the SoftSRV parameters.
In this section, we introduce the general SoftSRV framework along with a few specific parameterizations. First, we start with some basic notation and terminology.
Given a vocabulary $\mathcal{V}$ of textual tokens, let $\{x_{1},\ldots,x_{n}\}$ denote a sample of $n$ text sequences, belonging to the set of all possible sequences $S^{m}$ of a maximum length $m$, drawn according to a fixed but unknown distribution $\mathcal{D}$. Although we are not able to directly sample additional sequences from $\mathcal{D}$, our goal is to synthesize sequences that could have plausibly been drawn according to $\mathcal{D}$. We assume access to a (frozen) LLM, denoted $L:S^{m}\to S^{m}$, where we input and output sequences of equal fixed length $m$ for notational simplicity and without loss of generality. Furthermore, we explicitly decompose the LLM, $L=H\circ I$, where $I:S^{m}\to\mathbb{R}^{d\times m}$ represents the initial embedding layer that embeds each token of the input sequence to a $d$-dimensional dense vector, and $H:\mathbb{R}^{d\times m}\to S^{m}$ represents the remainder of the language model that maps the embedded tokens to the output sequence. For SoftSRV, we rely solely on the frozen model $H$ and we discard $I$ entirely. As explained below, we will replace $I$ by a parametrized family of embeddings.
The SoftSRV framework seeks to synthesize sequences similar to those drawn from $\mathcal{D}$ by training a dense embedding (or parameterized family of embeddings) $\mathbf{P}\in\mathbb{R}^{d\times t}$, with dimension $0<t<m$. We hereafter refer to $\mathbf{P}$ as the SoftSRV embedding. A successfully trained SoftSRV embedding, $\mathbf{P}$, should generate a sequence $x=H(\mathbf{P})$, via the frozen model $H$, that has a high likelihood of occurring under the distribution $\mathcal{D}$. More generally, we can sample several different sequences from a learned SoftSRV embedding, $x,x^{\prime},x^{\prime\prime}\ldots\sim H(\mathbf{P})$, by using randomized temperature-based decoding.
It is a recognized problem within the field of synthetic data generation that generated datasets often lack diversity (Li et al., 2023; Ben Allal et al., 2024). Although temperature sampling alone does allow for some variability, it is important to be able to inject more diversity in the generated sample. To this end, we further increase the variety of generated text by introducing a contextual SoftSRV embedding, $\mathbf{P}(\cdot):\mathbb{R}^{d_{e}}\to\mathbb{R}^{d\times t}$. A contextual SoftSRV embedding can be conditioned with different context vectors $\mathbf{z}\in\mathbb{R}^{d_{e}}$, during training and generation, to induce greater variations in the SoftSRV embedding.
@@
Before introducing specific parametrized families of SoftSRV embeddings, we describe the SoftSRV training procedure which is common throughout and also illustrated in Figure 1. We first let $\theta$ denote the trainable parameters of the contextual SoftSRV embedding $\mathbf{P}_{\theta}(\cdot)$. In addition to the sample of data ($x_{1},\ldots,x_{n}$) and frozen LLM ($H$), we assume access to a frozen sequence embedding function $E(\cdot):S^{m}\to\mathbb{R}^{d_{e}}$ that will define our context vector $\mathbf{z}_{i}=E(x_{i})$. During training, each training sequence is mapped to the context vector $\mathbf{z}_{i}=E(x_{i})$ and used to generate a conditioned SoftSRV embedding $\mathbf{P}_{\theta}(\mathbf{z}_{i})$, which is fed into the frozen LLM $H$ to produce a new sequence $x_{i}^{\prime}\sim H(\mathbf{P}_{\theta}(\mathbf{z}_{i}))$ using autoregressive next token generation. A standard causal (next-word) prediction loss, denoted $\ell(\cdot,\cdot)$, is backpropagated through the network up to the SoftSRV embedding layer $\mathbf{P}_{\theta}(\cdot)$, and an SGD-style update is applied to $\theta$ using the gradient $\nabla_{\theta}\ell(x_{i},x_{i}^{\prime})$. This loss can be thought of as a “reconstruction” error and the entire pipeline is akin to an auto-encoder. Viewing the pipeline through this lens, it is apparent that the sequence embedder $E(\cdot)$ should be sufficiently “lossy” in order to avoid making the learning problem trivial. This lossiness can be enforced by restricting the dimension $d_{e}$ of the embedding, for example.
@@
Once the contextual SoftSRV embedding $\mathbf{P}_{\theta}(\cdot)$ has been trained, we can then generate synthetic data by passing $\mathbf{P}_{\theta}(\mathbf{z})$ to the frozen LLM as embedded input context for different choices of context vector $\mathbf{z}$. A natural choice is to sample embeddings $(\mathbf{z}_{1},\ldots,\mathbf{z}_{n})$ derived from the data sample set $(x_{1},\ldots,x_{n})$.We now introduce a few specific SoftSRV parameterizations studied in this work.
2.1 SoftSRV Non-contextual Parameterization ($\mathrm{SS}_{\mathrm{NP}}$)
@@
The simplest parameterization treats the $dt$ entries of a SoftSRV embedding  $\mathbf{P}\in\mathbb{R}^{d\times t}$, directly as trainable parameters, i.e., $\theta=\mathbf{P}$ , resulting in the following objective:
$\mathrm{argmin}_{\theta}\sum_{i=1}^{n}\ell(H(\mathbf{P}),x_{i})\,\,.$ (1)
@@
This parameterization is an instance of a non-contextual SoftSRV embedding  i.e., any context $\mathbf{z}$ is ignored. Despite the lack of varying context, the synthesized output may still be diversified by using non-greedy (i.e., temperature sampling) decoding during LLM generation.
2.2 SoftSRV Mixture Parameterization ($\mathrm{SS}_{\mathrm{MPk}}$)
@@
Here, we train $k$ “basis” SoftSRV embedding matrices and define the final SoftSRV embedding  as a mixture of these bases. More precisely, in this variant the parameter set is $\theta=\{\mathbf{P}_{1},\ldots,\mathbf{P}_{k},\phi\}$, where $\mathbf{P}_{i}\in\mathbb{R}^{d\times t}$ are the basis SoftSRV embeddings
@@
and $W_{\phi}(\cdot):\mathbb{R}^{d_{e}}\to\mathbb{R}^{k}$ is a learned softmax function with parameters $\phi\in\mathbb{R}^{d_{w}}$. The trained $\mathrm{SS}_{\mathrm{MPk}}$ embedding is then the SGD solution to $\mathrm{argmin}_{\theta}\sum_{i=1}^{n}\ell(H(\mathbf{P}_{\theta}(E(x_{i})),x_{i})$.
The intuition behind this formulation is for each learned basis embedding $\mathbf{P}_{i}$ to encode a different aspect (mode) of the target data distribution and have each training example $x_{i}$ approximated by a mixture of these modes (similar to the intuition behind mixture or topic models (Hand, 2018)).
2.3 SoftSRV MLP Concatenated Parameterization ($\mathrm{SS}_{\mathrm{MC}}$)
Next, we consider a collection of $t$ small MLPs, whose output is concatenated to generate the final SoftSRV embedding. Let $F_{\phi_{i}}:\mathbb{R}^{d_{e}}\to\mathbb{R}^{t}$ denote the $i$th MLP with parameters $\phi_{i}$, and $\theta=\{\phi_{1},\ldots,\phi_{t}\}$ denote the parameters for the collection of MLPs. Then, we define:
@@ 2.3 SoftSRV MLP Concatenated Parameterization ($\mathrm{SS}_{\mathrm{MC}}$)
and the trained $\mathrm{SS}_{\mathrm{MC}}$ embedding is the SGD solution to $\mathrm{argmin}_{\theta}\sum_{i=1}^{n}\ell(H(\mathbf{P}_{\theta}(E(x_{i})),x_{i})$.
This parameterization is the most expressive that we consider, in that each embedding column (which can be thought of as an “soft token”) is computed using a distinct non-linear transformation of the context vector $\mathbf{z}$.
2.4 Differences with Prompt-tuning
Although SoftSRV and existing prompt-tuning methods (Lester et al., 2021; Li & Liang, 2021) both train embeddings, we outline several crucial differences between the two approaches. (1) The application: prompt-tuning is generally used to prepend a prompt embedding to an existing natural language prompt that is then fed to an LLM, with the goal of providing a higher quality response; in this work, our goal is to generate synthetic training data by feeding an LLM a trained embedding alone, i.e. there is no natural notion of an input/output pair in this application. (2) Contextual embedding: in standard prompt-tuning, it typically suffices to train a single prompt embedding that is prepended to different input prompts; in our setting, since we feed only the prompt embedding (and no natural language prompt) to the LLM we must do more than train a static embedding if we wish to induce significant variation. To that end, we introduce and train parameterized contextual embeddings. (3) The optimization procedure: in standard prompt-tuning the input/output sequence pairs (taken naturally from the application) can be used to train the prompt embedding; in the case of SoftSRV we train using an autoencoder-like compression and reconstruction scheme. For a discussion of the prompt-tuning literature, please see Appendix B.